OpenStudy (anonymous):

What are the solutions to the equation? x^2= 625? I need a couple of pointers. I'm a little lost

OpenStudy (anonymous):

Take the square root of both sides

OpenStudy (anonymous):

25 * 25 is 625

OpenStudy (anonymous):

Be careful square roots can be positive or negative.

OpenStudy (anonymous):

-625 + 625 + x2 = 0 + 625 -625 + 625 = 0 0 + x2 = 0 + 625 x2 = 0 + 625 x2 = 625 x2 = 625 x = {-25, 25}

OpenStudy (anonymous):

so there are two solutions -25 and 25

OpenStudy (anonymous):

I wish I could give all of you a Medal but Thank you so much! I now see how easy that was, and thank you Mihirb (:

OpenStudy (anonymous):

Your welcome :)

OpenStudy (radar):

\[x ^{2}-625=0\]\[(x+25)(x-25) = 0\]\[x=25\]\[x=-25\]